OpenStudy (anonymous):

How would you solve 23^x=6? I think you have to take log from both sides but I don't know what to do after that.

OpenStudy (anonymous):

What do you get when you do that?

OpenStudy (anonymous):

xlog23=log6

OpenStudy (anonymous):

Now try putting both sides over e, the log should cancel out, leaving (e^x)(23)=6

OpenStudy (anonymous):

Simplify to e^x=6/23

OpenStudy (anonymous):

That was the longer way around btw, didn't know if you knew how log/log worked

OpenStudy (anonymous):

Log/log is just change of base formula isnt it?

OpenStudy (anonymous):

Now take log of both sides, e will go away leaving x=log(6/23)=log6-log23

OpenStudy (anonymous):

Should be easier to find the answer now with a calculator

OpenStudy (anonymous):

I probably could have done that quicker still, but that's how I would have done it if I was still experimenting

OpenStudy (anonymous):

Would the answer then be x=log(6/23)?

OpenStudy (anonymous):

You could leave it like that too, I like the other form better imo =P

OpenStudy (anonymous):

Awesome! Thanks, I really struggled on the whole log concept, it's pretty complicated imo